OpenStudy (anonymous):

HELP WITH LOG FUNCTIONS! solve for x: 3log(x)=6-2x

OpenStudy (mertsj):

On the left side, use the property : mloga = loga^m

OpenStudy (anonymous):

ok so logx^3

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[\log(x^3)=6-2x\]

OpenStudy (anonymous):

thats the most you can simplify it?

OpenStudy (mertsj):

\[logx=2-2/3x\]

OpenStudy (mertsj):

I don't think you can solve this by ordinary algebraic methods. The best thing would be to graph both sides of the equation and see where they intersect.